OpenStudy (viola):

a bank is advertising that new customers can open a savings account with a 3 3/4 % interest rate compound annually. Robert invests $5000 in an account at this rate. if he makes no additional deposits or withdrawals on his account find the amount of money he will have to the nearest cent after 3 years?

OpenStudy (anonymous):

5000(1 + 0.0375)(1 + 0.0375)(1 + 0.0375) = 5000(1 + 0.0375)^3 = 5000(1.0375)^3 The "1" is from preserving the principal which gets added to the interest factor.

OpenStudy (anonymous):

All good now, @Viola ?

OpenStudy (viola):

Yes, thank you @tcarroll010, but can you explain in simpler words why I have to add "1"?

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Viola It's because after the first year, you have the principal: 5000 x 1 and the interest 5000 x 0.0375 so, 5000 x (1 + 0.0375) and repeat that for 2 more years

OpenStudy (viola):

I understand now, thank you for all your help! @tcarroll010

OpenStudy (anonymous):

If you did only 5000(0.0375) without the "1", it's like throwing away the principal

OpenStudy (anonymous):

uw!